# The non-Abelian geometry, topology, and dynamics of a nonreciprocal Su-Schrieffer-Heeger ladder

### v1

@@ Abstract
Non-Hermiticity breaks down the adiabaticity and naturally leads to the non-Abelian behaviors in multi-band systems. Here we consider a multi-band, non-Hermitian ladder model with the two legs being the nonreciprocal Su-Schrieffer-Heeger chains. We thoroughly study how the non-Abelian geometry, topology, and dynamics emerge in this model at the onset of inter-leg coupling. Under periodic boundary conditions, by defining a gauge-invariant winding number for chiral symmetric systems, we analytically give the exact topological phase diagram. With the aid of underlying symmetries generalized for non-Hermitian systems, we further refine the phase diagram by the geometry of band structure. In the pseudo-Hermitian symmetric regime, we find that the stable non-Abelian dynamics of a Bloch state under an external constant force can be well described in some conditions of the force by the Wilson line constructed for non-Hermitian systems. Under open boundary conditions, we also find that the bulk-boundary correspondence survives in the thermodynamic limit but breaks down for finite-size systems with the leg-dependent non-Hermitian skin effect (NHSE), demonstrating the so-called critical NHSE, of which the decaying length of the bulk skin modes $\xi$ varies with the system size $L$ and is numerically verified to satisfy the scale-free power law $\xi\propto L$. Our work may stimulate more focuses on the non-Abelian properties of the non-Hermitian/open quantum systems.
@@
In closed condensed matter systems, the geometry and topology of the band structure of a crystal is essential for understanding the electronic dynamics [1], where the key information comes from the Berry connection and the corresponding Berry phase of a Bloch state’s adiabatic evolution in momentum space (i.e., $k$-space) of the crystal [2], especially in single bands. The geometry and topology of the subspace involving multiple bands (called the non-Abelian geometry and topology in the following) can be well described by the non-Abelian Berry connection (i.e., Wilczek-Zee connection) and the corresponding Wilson lines [3] if the gaps to the other bands are very large. In this case, the inter-band evolution (called non-Abelian dynamics in the following) in the subspace is no longer adiabatic, although the overall evolution in the subspace preserves the adiabaticity to the subspaces of other bands. Non-Abelian properties are common in condensed matter systems, such as in topological insulators [4, 5] and graphenes [6], which cannot be properly characterized by single-band quantities due to the existence of band degeneracy in $k$-space. The non-Abelian topology and dynamics have been traced by the Wilson lines or in cold-atom experiments [7, 8] and been studied with various theoretic methods [9, 10, 11, 12].
@@
On the other hand, in open quantum systems the short-time dynamics can be effectively described by non-Hermitian Hamiltonians [13, 14]. The relaxation of the restriction of Hamiltonian’s Hermiticity brings about many unique phenomena to non-Hermitian systems, such as parity-time (PT) symmetry breaking [15, 16], non-unitary evolution, exceptional points [17], etc., among which the emergence of non-Hermitian skin effects (NHSEs) [18] is the underlying reason to break down and reconstruct the bulk-boundary correspondence (BBC) [19, 20, 21, 22, 23, 24, 25, 26, 27], and thus deepens our understanding of topology and dynamics in non-Hermitian systems [13].
@@
Since the intrinsic dissipation in non-Hermitian systems naturally breaks down the adiabaticity, leading to the mode switching between different bands, one may ask a question: what are the non-Abelian properties like in non-Hermitian systems? In Hermitian systems, one of the simplest theoretic models in one dimension with nontrivial non-Abelian properties is the two-leg Su-Schrieffer-Heeger (SSH) ladder [9] that couples two SSH chains as its two legs [28] and thus forms a four-band system. Different varieties of the two-leg SSH ladder are studied for broad theoretic interests [29, 30, 31, 32, 33, 34, 35, 36]. Because of the nonsymmorphic symmetry, the band spectrum of a Hermitian model always has a band crossing point [37, 38, 39, 40, 41], which invalidates the single-band approximation and requires the non-Abelian quantities to describe the topology and dynamics. Here, we generalize this two-leg SSH ladder to the non-Hermitian regime that instead couples two nonreciprocal SSH chains with opposite NHSEs. Apparently, this non-Hermitian model is a four-band system like the Hermitian counterpart but with a complex spectrum, and the nonreciprocal legs offer the nontrivial non-Hermitian topology as well as the NHSEs. Therefore, the non-Hermitian two-leg SSH ladder offers a simple non-Hermitian platform to investigate the non-Abelian properties of non-Hermitian systems. In addition, the coupling of two nonreciprocal SSH chains with different skin depths will lead to the so-called critical NHSE [42, 43, 44, 45]. In this non-Hermitian two-leg SSH ladder, one may also observe how the non-Abelian properties emerge from the inter-leg coupling.
@@
Although there have been several works on different aspects on non-Abelian properties of non-Hermitian systems, such as the quantum metric in pseudo-Hermitian symmetric phases [46], knots and Wilson lines of non-Hermitian bands [47], the PT symmetry breaking [48], non-Abelian Bloch dynamics [49], the gain/loss effect in spin-orbit coupled systems [50], and also some works related to other non-Hermitian two-leg ladders [51, 52, 53, 54, 55, 56, 57], no detailed analysis is found for the non-Abelian geometry, topology, and dynamics of the non-Hermitian two-leg SSH ladder with opposite NHSEs. Therefore, in this paper after introducing the effective non-Hermitian Hamiltonian from the Lindblad master equation in Sec. I, based on the symmetry analysis of the Hamiltonian in Sec. II, we analytically give the exact non-Abelian topological phase diagram with various geometries of band structure via a newly defined gauge-invariant winding number under periodic boundary conditions (PBCs) in Sec. III, numerically demonstrate and understand the non-Abelian dynamics in the pseudo-Hermitian symmetric regime under different strengths of the external constant force with the aid of non-Hermitian Wilson lines in Sec. IV, and verify the BBC under open boundary conditions (OBCs) in the thermodynamic limit and the critical NHSE for the finite-size systems with the scale-free power law after the onset of inter-leg coupling in Sec. V. Finally, Sec. VI summarizes the results.
I Effective Hamiltonian of the dissipative two-leg SSH ladder
@@
Figure 1: (a) Upper panel: The sketch of a dissipative two-leg (denoted by $\uparrow$ and $\downarrow$) SSH ladder coupled to the environment (denoted by the curly arrows), where $\kappa,\nu$, and $\chi$ label the corresponding reciprocal hoppings. Lower panel: The sketch of the effective non-Hermitian Hamiltonian $\hat{H}_{\text{NH}}$ formed by coupling two SSH chains with opposite nonreciprocal hoppings labeled by $\kappa_{\pm}$. The other parameters are the same as those in the upper panel. The gray dotted rectangle circles out the $j$th unit cell with $A$- and $B$-sublattice sites. (b) A typical topological phase diagram under PBCs characterized by the winding number $w=1$ (blue), $-1$ (yellow), and 00 (white), with different geometries of band structures characterized by real line gaps (gray vertical lines), imaginary line gaps (gray horizontal lines), and point gaps (gray dots). The black solid lines indicate the gap closing/phase transition. The purple dashed line separates the purely real spectra from the others. The red vertical line at $\chi^{2}=0$ separates the negative ($-\chi^{2}$) and positive $(\chi^{2})$ cases and also indicates the singularity caused by the critical NHSE under OBCs. The orange dotted line at $\chi^{2}=4$ is the parameter path for Fig.3(a). (c-r) Typical geometries of band structures for the points denoted by red triangles in (b) with parameters $(\chi^{2},\delta^{2})=(0.8,13.5)$, $(5,13)$, $(11.5,13)$, $(5,10)$, $(10,10)$, $(2,7.5)$, $(2,7.5)$, $(5,7)$, $(5,6.2)$, $(10,5)$, $(3,3)$, $(3.5,1.5)$, $(0.4,2)$, $(0.5,1)$, $(3.5,0.3)$, and $(1,0.2)$, respectively. Bottom gray layers: the projection of the spectra under PBCs on the complex energy plane; Top gray layers: the corresponding spectra under OBCs for $L=200$, where the isolated dots at zero energy in some gaps are the end states in the topological regimes.
@@
Here we couple the system, a two-leg SSH ladder [9] that is formed by two SSH chains relatively shifted by half unit cell, to the environment, as shown in the upper panel of Fig. 1(a). Under the Markov approximation, the dynamics can be well described by the Lindblad master equation [14] as follows:
@@
represents a two-leg SSH ladder with $\hat{s}^{({\dagger})}_{j\sigma}$ ($s=a,b$ and $\sigma=\,\uparrow,\downarrow$) being the annihilation (creation) operator at the $s$-sublattice site of the $j$th unit cell in leg-$\sigma$ of the ladder, and system parameters $\{\kappa,\nu,\chi_{\pm}\}$, which are set real for simplicity, being the hopping strengths shown in Fig. 1(a). The jump operators are set as [58]
@@
$\frac{d\hat{\rho}}{dt}=-i\big{(}\hat{H}_{\text{NH}}\hat{\rho}-\hat{\rho}\hat{H}^{\dagger}_{\text{NH}}\big{)}+2\sum_{j,\sigma}\hat{L}_{j\sigma}\hat{\rho}\hat{L}_{j\sigma}^{{\dagger}},$ (4)
where the summation term is called the quantum jump term. One can use the effective non-Hermitian Hamiltonian, $\hat{H}_{\text{eff}}\equiv\hat{H}-i\sum_{j\sigma}{\hat{L}_{j\sigma}^{{\dagger}}\hat{L}_{j\sigma}}$, to capture the short-time dynamics before a quantum jump occurs, or can directly neglect the quantum jump term by the schemes of post-selection in experiments. In the explicit form, one can easily find that the effective Hamiltonian,
@@
$+\,\hat{b}_{j\downarrow}^{{\dagger}}\hat{a}_{j+1,\downarrow})+[\nu(\hat{b}_{j\uparrow}^{\dagger}\hat{a}_{j+1,\uparrow}+\hat{a}_{j\downarrow}^{\dagger}\hat{b}_{j\downarrow})+\chi_{+}\hat{a}_{j\uparrow}^{{\dagger}}\hat{a}_{j\downarrow}$
$+\,\chi_{-}\hat{b}_{j\uparrow}^{{\dagger}}\hat{b}_{j\downarrow}+\text{H.c.}]\big{\}}-i\delta\sum_{j\sigma}(\hat{a}^{\dagger}_{j\sigma}\hat{a}_{j\sigma}+\hat{b}^{\dagger}_{j\sigma}\hat{b}_{j\sigma})$
@@
is just a non-Hermitian generalization of the two-leg SSH ladder by including a nonreciprocal hopping denoted by $\kappa_{\pm}=\kappa\pm\delta$ and an overall loss denoted by the total particle number operator $\hat{N}$. Note that in our case, the nonreciprocal hoppings in two legs are reversed such that opposite NHSEs occur for two legs in the limit of vanishing inter-leg coupling.
Since the loss term only contributes an overall decaying factor to the dynamics, without loss of generality, in the following we just focus on the Hamiltonian $\hat{H}_{\text{NH}}$ instead, as shown in the lower panel of Fig. 1(a).
II Symmetries of the complex band structure
To obtain the spectrum in $k$-space, we take the discrete Fourier transform,
@@ II Symmetries of the complex band structure
where $L$ is the number of unit cells, $d$ is the length of the unit cell, and the summation is done over one Brillouin zone (BZ) with $k=2\pi n/L$ ($n\in$ Integer). In the following, we set $d=\nu=1$ as the units of length and energy, respectively. The non-Hermitian Hamiltonian in Eq. (5) can be transformed from the real space to $k$-space under PBCs:
@@
where the bases in both spaces respectively read $\hat{\psi}^{\dagger}\equiv(\{\hat{a}^{\dagger}_{j\uparrow}~{}\hat{a}^{\dagger}_{j\downarrow}~{}\hat{b}^{\dagger}_{j\uparrow}~{}\hat{b}^{\dagger}_{j\downarrow}\})$ and $\hat{\psi}_{k}^{\dagger}=\big{(}\hat{a}^{\dagger}_{k\uparrow}~{}\hat{a}^{\dagger}_{k\downarrow}~{}\hat{b}^{\dagger}_{k\uparrow}~{}\hat{b}^{\dagger}_{k\downarrow}\big{)}$, and the Hamiltonian matrix in $k$-space is
$H_{k}$ $=$ $\left(\begin{matrix}\chi_{+}&0\\ 0&\chi_{-}\end{matrix}\right)\otimes\sigma_{x}+\tau_{+}\otimes\left(\begin{matrix}\kappa_{+}+e^{-ik}&0\\ 0&1+\kappa_{+}e^{-ik}\end{matrix}\right)$ (8)
$~{}+\tau_{-}\otimes\left(\begin{matrix}\kappa_{-}+e^{ik}&0\\ 0&1+\kappa_{-}e^{ik}\end{matrix}\right).$
@@
Hereafter, $\tau_{x,y,z}$ and $\sigma_{x,y,z}$ are Pauli matrices acting on the sublattice ($a,b$) and the leg ($\uparrow$,$\downarrow$ , ↓) spaces, respectively; $\tau_{0}$ and $\sigma_{0}$ are corresponding identity matrices, and $\tau_{\pm}\equiv(\tau_{x}\pm i\tau_{y})/2$.
@@
When $\chi_{+}=\chi_{-}$, due to the underlying symmetries of $\hat{H}_{\text{HN}}$ in real space (the details can be referred to in Appendix B), the Hamiltonian matrix $H_{k}$ in $k$-space satisfies the following relations inherited from the corresponding symmetries: (a) Time-reversal symmetry: $KH_{k}K=H_{k}^{*}=H_{-k}$ with $K$ being the complex conjugation, due to the reality of all parameters of $\hat{H}_{\text{NH}}$; (b) Sublattice symmetry: $CH_{k}C^{\dagger}=-H_{k}$ with $C\equiv\tau_{z}\otimes\sigma_{z}=C^{\dagger}=C^{-1}$, due to the bipartition of the ladder; (c) Pseudo-mirror symmetry: $MH_{k}M=H^{\dagger}_{-k}$ with $M\equiv\tau_{x}\otimes\sigma_{0}=M^{\dagger}=M^{-1}$, which is a non-Hermitian generalization of the traditional mirror symmetry; (d) Inversion symmetry: $I_{k}H_{k}I_{k}^{\dagger}=H_{-k}$ with $I_{k}=\left(\begin{array}[]{cc}1&0\\ 0&e^{-ik}\end{array}\right)\otimes\sigma_{x}$ satisfying $I_{k}^{\dagger}=I_{k}^{-1}$; (d) Pseudo-glide symmetry: $G_{k}H_{k}G_{k}^{\dagger}=H_{k}^{\dagger}$ with $G_{k}=\left(\begin{array}[]{cc}0&e^{-ik}\\ 1&0\end{array}\right)\otimes\sigma_{x}$ satisfying $G_{k}^{\dagger}=G_{k}^{-1}$ and $G_{k}^{2}=e^{-ik}I$ ($I$: identity matrix). These symmetries ensure that the eigenenergies of $H_{k}$ belong to the $D_{2h}$ point group in the composite 3D space spanned by the three 2-fold rotating axes of the $k$-axis, the real and the imaginary axes of the complex energy plane, which implies that the gap closing (indicating a phase transition) can only occur at the real or imaginary axes. The typical geometries of band structures are shown in Figs. 1(c)-1(r).
@@
Additionaly, besides the symmetries mentioned above, $H_{k}$ also has an $\eta_{k}$-pseudo-Hermitian symmetry (the recap of pseudo-Hermiticity can be found in Appendix A) [13]:
$\eta_{k}H_{k}\eta_{k}^{-1}=H_{k}^{\dagger}$ (9)
with
@@
$\eta_{k}\equiv e^{ik/2}G_{k}=\left(\begin{array}[]{cc}0&e^{-ik/2}\\ e^{ik/2}&0\end{array}\right)\otimes\sigma_{x}=\eta_{k}^{-1}=\eta_{k}^{\dagger}.$ (10)

### v2

# Non-Abelian geometry, topology, and dynamics of a nonreciprocal Su-Schrieffer-Heeger ladder

@@ Abstract
Non-Hermiticity naturally breaks down the adiabaticity and thus leads to non-Abelian behaviors in multi-band systems. Here, we study how non-Abelian properties emerge in non-Hermitian systems by considering a multi-band non-Hermitian model — the nonreciprocal Su-Schrieffer-Heeger (SSH) ladder that is formed by coupling two nonreciprocal SSH chains. Under periodic boundary conditions, we analytically obtain the exact phase diagrams of the geometry of band structure classified by its complex value and gap type, and of the non-Abelian topology based on a newly defined gauge-invariant winding number under the chiral symmetry. Under open boundary conditions, we find that the bulk-boundary correspondence survives in the thermodynamic limit but breaks down for finite sizes along with the emergence of critical non-Hermitian skin effects when the inter-leg coupling is weak, where the decaying length $\xi$ of the bulk skin modes varies with the system size $L$, satisfying the scale-free power law $\xi\propto L$. Finally, we demonstrate the non-Abelian dynamics of a Bloch state subject to an external constant force in the pseudo-Hermitian symmetric regime in comparison with the non-Hermitian Wilson lines. Our work may stimulate further interests in nontrivial non-Abelian behaviors in non-Hermitian and open quantum systems.
I Introduction
@@
In closed condensed matter systems, the geometry and topology of the band structure of a crystal are essential for understanding the electronic dynamics [1], where the key information comes from the Berry connection and the corresponding Berry phase of a Bloch state’s adiabatic evolution in momentum space (i.e., $k$ space) of the crystal [2], especially in single bands. The geometry and topology of the subspace involving multiple bands (called the non-Abelian geometry and topology) can be well described by the non-Abelian Berry connection (i.e., Wilczek-Zee connection) and the corresponding Wilson lines [3] if the gaps to the other bands are very large. In this case, the inter-band evolution (called the non-Abelian dynamics) in the subspace is no longer adiabatic, although the overall evolution in the subspace preserves the adiabaticity against the subspaces of other bands. Non-Abelian properties are common in condensed matter systems, such as in topological insulators [4, 5] and graphenes [6], which cannot be properly characterized by single-band quantities due to the existence of band degeneracy in $k$ space. The non-Abelian topology and dynamics can be traced by Wilson lines in cold-atom experiments [7, 8] and have been broadly studied with various theoretic methods [9, 10, 11, 12].
@@
On the other hand, in open quantum systems the short-time dynamics can be effectively described by non-Hermitian Hamiltonians [13, 14]. The relaxation of the Hamiltonian’s Hermiticity brings about many unique phenomena to non-Hermitian systems, such as parity-time-reversal (PT) symmetry breaking [15, 16], non-unitary evolution, exceptional points [17], etc., among which the emergence of the so-called non-Hermitian skin effects (NHSEs) [18] is one important mechanism for the breakdown and thus reconstruction of the bulk-boundary correspondence (BBC) [19, 20, 21, 22, 23, 24, 25, 26, 27], deepening our understanding of topology and dynamics in non-Hermitian systems [13].
@@
Since the intrinsic dissipation in non-Hermitian systems naturally breaks down the adiabaticity, leading to the repopulation at different bands, one may ask a question: What are the non-Abelian properties like in non-Hermitian systems? In Hermitian systems, one of the simplest models with nontrivial non-Abelian properties is the two-leg Su-Schrieffer-Heeger (SSH) ladder [9] that is formed by coupling two SSH chains as its two legs [28] and thus has four bands. Different varieties of the SSH ladder have been theoretically studied for broad interests [29, 30, 31, 32, 33, 34, 35, 36]. Because of the nonsymmorphic symmetry, the band structure of a Hermitian SSH ladder always supports crossing points in $k$ space [37, 38, 39, 40, 41], which invalidate the single-band description and require the non-Abelian quantities to describe the topology and dynamics. Here, we generalize this Hermitian SSH ladder to the non-Hermitian regime that instead couples two nonreciprocal SSH chains as its two legs, dubbed the nonreciprocal SSH ladder. Apparently, this non-Hermitian ladder is also a four-band model like its Hermitian counterpart but with a generally complex spectrum, and the nonreciprocal legs imply the nontrivial non-Hermitian topology regarding NHSEs. Therefore, the nonreciprocal SSH ladder offers a simple non-Hermitian platform to investigate the non-Abelian properties of non-Hermitian systems. In addition, the coupling of two nonreciprocal SSH chains with different skin depths will lead to the so-called critical NHSE [42, 43, 44, 45]. In this nonreciprocal SSH ladder, one may also observe how the non-Abelian properties emerge from the inter-leg coupling.
@@
Although several aspects of non-Abelian properties in non-Hermitian systems have been studied, such as non-Abelian Berry connections in pseudo-Hermitian symmetric phases [46], knots of non-Hermitian Bloch bands [47], PT symmetry breaking in non-Hermitian non-Abelian lattice models [48], non-Abelian dynamics in dissipative photonic lattices [49], and the gain and loss effect in spin-orbit coupled systems [50], and also there are some works regarding different versions of non-Hermitian two-leg ladders [51, 52, 53, 54, 55, 56, 57, 58], no detailed analysis is found for the non-Abelian geometry, topology, and dynamics of the nonreciprocal SSH ladder. Therefore, in this paper after introducing the nonreciprocal SSH ladder in Sec. II, under periodic boundary conditions (PBCs) we analytically obtain the exact phase diagrams of the geometry of band structure classified by its complex value and gap type in Sec. III, and of the non-Abelian topology based on a newly-defined gauge-invariant winding number in Sec. IV, where the BBC in the thermodynamic limit and the critical NHSE with the scale-free power law for finite sizes under open boundary conditions (OBCs) are also demonstrated. In Sec. V, we study the non-Abelian dynamics of a Bloch state in the pseudo-Hermitian symmetric regime subject to an external constant force in comparison with the non-Hermitian Wilson lines. Finally, Sec. VI gives a conclusion and a short discussion on experiments.
II The nonreciprocal SSH ladder
II.1 Effective non-Hermitian Hamiltonian
@@
Figure 1: (a) Upper panel: Sketch of a dissipative two-leg (denoted by $\uparrow$ and $\downarrow$) SSH ladder coupled to the environment (denoted by the curly arrows), where $\kappa,\nu$, and $\chi_{\pm}$ label the corresponding hoppings. Lower panel: Sketch of the nonreciprocal SSH ladder represented by $\hat{H}_{\text{NH}}$ with $\kappa_{\pm}$ labeling the nonreciprocal hoppings and other parameters being the same as in the upper panel. The rectangles circle out the unit cells including AA- and $B$-sublattice sites. (b) A typical phase diagram under PBCs for $\kappa=2.5$, composed of the geometry of band structure, characterized by real line gaps (gray vertical lines), imaginary line gaps (gray horizontal lines), and point gaps (gray dots), and of the non-Abelian topology, characterized by the winding number $w=1$ (blue), 1-1 (yellow), and 0 (white). The black solid lines are the gap-closing positions. Below the purple dashed curves are the pseudo-Hermitian symmetric regimes. The red vertical line at $\chi=0$ separating the negative ($\chi^{2}$) and positive $(\chi^{2})$ cases are topologically singular in the thermodynamic limit. The orange dotted line at $\chi^{2}=4$ is the parameter path for Fig. 2(a). (c)–(r) Typical band structures in the composite 3D space spanned by $k$ and the complex energy plane for parameters $(\chi^{2},\delta^{2})=(0.8,13.5)$, $(5,13)$, $(11.5,13)$, $(5,10)$, $(10,10)$, $(2,7.5)$, $(2,7.5)$, $(5,7)$, $(5,6.2)$, $(10,5)$, $(3,3)$, $(3.5,1.5)$, $(0.4,2)$, $(0.5,1)$, $(3.5,0.3)$, and $(1,0.2)$, respectively, labeled by red triangles in (b). Bottom gray layers: the projection of band structures under PBCs onto the complex energy plane; Top gray layers: the corresponding energy spectra under OBCs for $L=200$, where the isolated dots in some gaps are the topological zero-mode end states.
@@
Here we couple the system, a two-leg SSH ladder [9] formed by two SSH chains that are relatively shifted by half unit cell, to the environment, as shown in the upper panel of Fig. 1(a). Under the Markov approximation, the dynamics can be well described by the Lindblad master equation [14] as follows:
@@
$+\big{(}\chi_{+}\hat{a}_{j\uparrow}^{{\dagger}}\hat{a}_{j\downarrow}+\chi_{-}\hat{b}_{j\uparrow}^{{\dagger}}\hat{b}_{j\downarrow}\big{)}+\text{H.c.}\big{]},$
@@
$\frac{d\hat{\rho}}{dt}=-i\big{(}\hat{H}_{\text{eff}}\hat{\rho}-\hat{\rho}\hat{H}^{\dagger}_{\text{eff}}\big{)}+2\sum_{j,\sigma}\hat{L}_{j\sigma}\hat{\rho}\hat{L}_{j\sigma}^{{\dagger}},$ (4)
where the summation term is the quantum jump term, and the other term with the effective Hamiltonian, $\hat{H}_{\text{eff}}\equiv\hat{H}-i\sum_{j\sigma}{\hat{L}_{j\sigma}^{{\dagger}}\hat{L}_{j\sigma}}$, can capture the short-time dynamics before a quantum jump occurs with the aid of post-selection techniques in experiments. In the explicit form, one can find that the effective Hamiltonian,
@@
$+\,\hat{b}_{j\downarrow}^{{\dagger}}\hat{a}_{j+1,\downarrow})+\big{[}\nu(\hat{b}_{j\uparrow}^{\dagger}\hat{a}_{j+1,\uparrow}+\hat{a}_{j\downarrow}^{\dagger}\hat{b}_{j\downarrow})+\chi_{+}\hat{a}_{j\uparrow}^{{\dagger}}\hat{a}_{j\downarrow}$
$+\,\chi_{-}\hat{b}_{j\uparrow}^{{\dagger}}\hat{b}_{j\downarrow}+\text{H.c.}\big{]}\Big{\}}-i\delta\sum_{j\sigma}(\hat{a}^{\dagger}_{j\sigma}\hat{a}_{j\sigma}+\hat{b}^{\dagger}_{j\sigma}\hat{b}_{j\sigma})$
@@
is just a non-Hermitian generalization of the two-leg SSH ladder by introducing nonreciprocal hoppings denoted by $\kappa_{\pm}=\kappa\pm\delta$ and an overall loss related to the total particle number operator $\hat{N}$. Note that in this model, the nonreciprocal hoppings along two legs are engineered to be relatively reversed such that reversed NHSEs occur in different legs when the inter-leg coupling vanishes.
Since the loss term only contributes an overall decaying factor to the dynamics, without loss of generality, in the following we instead focus on the non-Hermitian Hamiltonian $\hat{H}_{\text{NH}}$ of the nonreciprocal SSH ladder that is formed by coupling two nonreciprocal SSH chains, as shown in the lower panel of Fig. 1(a).
II.2 Hamiltonian in $k$ space and the symmetries
To obtain the non-Abelian properties in $k$ space, we transform the non-Hermitian Hamiltonian $\hat{H}_{\rm NH}$ in Eq. (5) from real space to $k$ space under PBCs:
via the discrete Fourier transform,
@@ II.2 Hamiltonian in $k$ space and the symmetries
where $L$ is the number of unit cells, $d$ is the length of the unit cell, and the summation is done over one Brillouin zone (BZ) with $k=2\pi n/L$ ($n\in$ Integer). The bases in both spaces, respectively, read as $\hat{\psi}^{\dagger}\equiv(\{\hat{a}^{\dagger}_{j\uparrow}~\hat{a}^{\dagger}_{j\downarrow}~\hat{b}^{\dagger}_{j\uparrow}~\hat{b}^{\dagger}_{j\downarrow}\})$ and $\hat{\psi}_{k}^{\dagger}=\big{(}\hat{a}^{\dagger}_{k\uparrow}~\hat{a}^{\dagger}_{k\downarrow}~\hat{b}^{\dagger}_{k\uparrow}~\hat{b}^{\dagger}_{k\downarrow}\big{)}$, and the Hamiltonian matrix in $k$ space
$H_{k}$ $=$ $~\left(\begin{matrix}\chi_{+}&0\\ 0&\chi_{-}\end{matrix}\right)\otimes\sigma_{x}$ (8)
$~+\tau_{+}\otimes\left(\begin{matrix}\kappa_{+}+\nu e^{-ikd}&0\\ 0&\nu+\kappa_{+}e^{-ikd}\end{matrix}\right)$
$~+\tau_{-}\otimes\left(\begin{matrix}\kappa_{-}+\nu e^{ik}&0\\ 0&\nu+\kappa_{-}e^{ikd}\end{matrix}\right),$
@@
where $\tau_{x,y,z}$ and $\sigma_{x,y,z}$ are Pauli matrices acting on the sublattice ($a,b$) and the leg ($\uparrow,\downarrow$) spaces, respectively; $\tau_{0}$ and $\sigma_{0}$ are corresponding identity matrices, and $\tau_{\pm}\equiv(\tau_{x}\pm i\tau_{y})/2$. In the following, we set $d=\nu=1$ as the units of length and energy, respectively.
@@
The symmetries are useful tools for analyzing non-Abelian properties. Due to the underlying symmetries of $\hat{H}_{\text{NH}}$ in real space, $H_{k}$ satisfies the following relations inherited from the corresponding symmetries: (a) time-reversal symmetry: $H_{k}^{*}=H_{-k}$, due to the reality of all parameters in $\hat{H}_{\text{NH}}$; (b) chiral symmetry: $CH_{k}C^{\dagger}=-H_{k}$ with $C\equiv\tau_{z}\otimes\sigma_{z}=C^{\dagger}=C^{-1}$, due to the bipartition of the ladder; (c) inversion symmetry: $I_{k}H_{k}I_{k}^{\dagger}=H_{-k}$ with $I_{k}=\left(\begin{array}[]{cc}1&0\\ 0&e^{-ik}\end{array}\right)\otimes\sigma_{x}$ satisfying $I_{k}^{\dagger}=I_{k}^{-1}$. These relations ensure that the band structure of $H_{k}$ has a $D_{2h}$ point group symmetry in the composite three-dimensional (3D) space spanned by three twofold rotating axes: the $k$ axis, the real and the imaginary axes of the complex energy plane. This implies that the gap closing (indicating a phase transition) can only occur at the real or imaginary axes, as shown in Figs. 1(c)–1(r). The details of symmetry analysis can be referred to in Appendix B.
@@
Additionally, $H_{k}$ also has an $\eta_{k}$-pseudo-Hermitian symmetry (the recap of pseudo-Hermiticity can be found in Appendix A) [13]:
$\eta_{k}H_{k}\eta_{k}^{-1}=H_{k}^{\dagger},$ (9)
where the invertible Hermitian matrix $\eta_{k}$ can be explicitly expressed for $\chi_{+}=\chi_{-}$ as
@@
$\eta_{k}\equiv\left(\begin{array}[]{cc}0&e^{-ik/2}\\ e^{ik/2}&0\end{array}\right)\otimes\sigma_{x}=\eta_{k}^{-1}=\eta_{k}^{\dagger},$ (10)